OpenStudy (anonymous):

A bag contains 3 black marbles and 2 white marbles. Jake reaches in and selects one marble without looking. What is the probability that he gets a white marble? twenty percent forty percent fifty percent sixty-seven percent

OpenStudy (johnweldon1993):

How many possible outcomes are there? Well that would be ALL the marbles...so 3 + 2 = 5 Out of those 5 possible outcomes...we want to pick a white marble...there are 2 of those... So our probability would be \[\large \frac{\text{Outcomes we want}}{\text{Possible outcomes}} = \frac{2}{5} = \space ?\]